# It’s Not Easy Being Green: On the Energy Efficiency of Programming Languages

### v2

@@ 3.2.2. Performance counters
Section 4.4.2 highlights the impact of memory activity on energy consumption: a majority of memory energy consumption is dependent on the rate of read and write operations. We monitor the last level cache (LLC) using performance counters to estimate memory activity. Each LLC hit means that the cache already had a copy of the data, and therefore no further memory activity occurs. On the other hand, each LLC miss means that the system must fetch data from memory. Therefore, the number of LLC misses is a good proxy for memory activity.
@@
3.3. Benchmarks
This paper primarily focuses its evaluation on the benchmarks used in the original Pereira et al. study, namely the Computer Language Benchmark Game. In some analyses, we extend that evaluation with two additional benchmark suites: SPEC CPU 2017 (Corporation, 2022) and DaCapo Chopin (Blackburn et al., 2025). The following subsections provide a brief overview of each of these benchmark suites.
3.3.1. Computer Language Benchmark Game
Table 3. Descriptions of benchmarks from the Computer Language Benchmark Game (§3.3.1), along with the range of lines of code (LoC) across languages for each benchmark as reported by cloc.
@@ 3.3.1. Computer Language Benchmark Game
binary-trees 36–98 Allocates and frees binary trees of various specified depth.
fannkuch-redux 40–158 Computes the max number of flips to sort a permutation.
fasta 84–287 Generates random DNA sequences.
k-nucleotide 57–202 Counts frequencies of nucleotide sequences in a string.
mandelbrot 32–140 Plots the Mandelbrot set $\left[-1.5-i,0.5+i\right]$.
n-body 78–157 Models orbits of Jovian planets.
pidigits 41–149 Generates digits of $\pi$ using arbitrary precision arithmetic.
regex-redux 22–111 Uses simple regex patterns to manipulate DNA data.
reverse-complement 34–257 Outputs complements of each nucleotide from sequence.
spectral-norm 33–156 Calculates the spectral norm of an infinite matrix.
The Computer Language Benchmark Game (Gouy, 2024) (CLBG) is a collection of small programs implemented in many different programming languages (22 to 287 lines of code). We use the same benchmark implementations as Pereira et al. (Couto and Pereira, 2017), which is in effect a snapshot of the fastest versions of the benchmarks as available on the CLBG repository at the time of the original work. Table 3 provides a brief description of each benchmark. We limit our analysis to eleven languages across thirteen implementations: C, C#, C++, Go, Java, JavaScript, Lua (Lua and LuaJIT), PHP, Python (CPython and PyPy), Rust, and TypeScript. These languages comprise 11 of the top 12 most popular general programming languages in the 2024 StackOverflow survey (Overflow, 2024) (excluding Kotlin).
@@
We stress that while the CLBG benchmarks themselves are not necessarily representative of real-world applications, the causal analysis this paper develops is largely independent of the details of the benchmark implementations. It instead highlights the impact of high-level properties of the benchmark implementations, such as their degree of parallelism and cache activity.
3.3.2. SPEC CPU 2017
Table 4. SPEC CPU 2017 benchmarks (§3.3.2).
@@ 3.3.2. SPEC CPU 2017
600.perlbench_s C Perl interpreter
602.gcc_s C GNU C compiler
605.mcf_s C Route planning
620.omnetpp_s C++ Discrete-event simulation
623.xalancbmk_s C++ XML to HTML conversion
625.x264_s C Video compression
631.deepsjeng_s C++ Alpha-beta tree search
641.leela_s C++ Monte Carlo tree search
648.exchange2_s Fortran Recursive solution generator
657.xz_s C General data compression
603.bwaves_s Fortran Explosion modeling
607.cactuBSSN_s C++, C, Fortran Physics: relativity
619.lbm_s C Fluid dynamics
621.wrf_s Fortran, C Weather forecasting
627.cam4_s Fortran, C Atmosphere modeling
628.pop2_s Fortran, C Wide-scale ocean modeling
638.imagick_s C Image manipulation
644.nab_s C Molecular dynamics
649.fotonik3d_s Fortran Computational electromagnetics
654.roms_s Fortran Regional ocean modeling
Table 5. DaCapo benchmarks (§3.3.3).
@@
SPEC CPU 2017 (Corporation, 2022) is a widely used benchmark suite for evaluating systems on C, C++ and Fortran code. The suite consists of large real-world programs and workloads, ranging between 1k and 1.3M lines of code. We use the latest available version (1.1.9) and run all benchmarks in the “speed” category using the reference input set. We omit a single benchmark, 627.cam4_s, as it did not successfully run on our platform. Table 4 provides a brief description of each benchmark.
3.3.3. DaCapo
DaCapo (Blackburn et al., 2025) is a widely used benchmark suite for Java applications. The suite consists of large real-world programs and workloads, ranging between 25k and 6M lines of code. We use the latest available version (Chopin) and run all benchmarks using the default input set. Table 5 provides a brief description of each benchmark.
4. Causal Model for Energy Consumption in Programming Languages
This section builds piece by piece the causal model presented as a diagram in its final form in Figure 2. Causal diagrams illustrate how different factors influence each other, with arrows indicating cause-to-effect relationships (Pearl, 2009). We additionally use gray arrows to represent comparatively weaker relationships, as detailed in the following sections. Each subsection below gradually adds nodes and edges to the diagram in order to obtain a rich causal model that captures the essential factors in the relationship between the choice of programming language and energy consumption. Figures 4, 5, 9, 11, and 13 build upon each other, with new elements in orange.
4.1. Starting Point
Figure 4. The simple model used as a starting point in this paper (§4.1). This simple model captures the relationship implied in Pereira et al., namely that the choice of programming language has a direct impact on total energy consumption.
[…]
The start of our working causal diagram with two nodes, “Programming Language” and “Energy”. There is an edge from “Programming Language” to “Energy”.
@@ 4.1. Starting Point
4.2. Programming Languages and Implementations
Figure 5. Programming languages may have multiple implementations (§4.2), and implementation decisions such as garbage collection or using just-in-time compilation have a more direct impact on performance. We are in fact comparing implementations of programming languages, not the languages themselves.
[…]
Our previous working causal diagram, adding “Programming Language Implementation”.
Figure 6. Execution time for the ten longest-running pyperformance benchmarks in the five last Python releases, normalized to CPython 3.9 (§4.2). The thicker gray line indicates $y=1$. Most benchmarks show improvements over time, with a $40\%$ speedup on average for the entire suite between CPython 3.9 and 3.13.
[…]
A line graph with 10 lines, one for each benchmark. 9 out of 10 benchmarks end under the $y=1$ line for CPython 3.13.
Figure 5 introduces the important distinction between the programming languages themselves and their implementations. Some languages have a single well-known implementation, or a blurry line between language and implementation, such as Rust or Go. Others have multiple implementations, such as C/C++ with GCC, Clang, or MSVC, Java with various JVMs, Python with CPython and PyPy, or Lua with LuaJIT. Finally, there can be significant differences between specific versions of the same implementation. Our model treats those different versions as different implementations altogether. For example, we test the pyperformance benchmark suite (pyperformance, 2025) across CPython versions 3.9 through 3.13. Figure 6 shows overall improvements for the ten longest-running benchmarks. Over the entire suite, we observe a $40\%$ speedup using CPython 3.13 over 3.9.
@@ 4.2. Programming Languages and Implementations
Figure 7. Programming Languages can exhibit different performance characteristics depending on the implementation used (§4.2). Here, we compare interpreters (CPython, Lua) to their JIT equivalent (PyPy, LuaJIT) on the same benchmark sources. We normalize to the interpreter results. We find around 5×5$\times$5 × and 1.25×1.25$\times$1.25 × decreased execution time and energy consumption on average for LuaJIT over Lua and PyPy over CPython, respectively. These results demonstrate the impact of choice of programming language implementation versus the language itself.
[…]
Two bar charts illustrating relative execution time and energy consumption between CPython/PyPy and Lua/LuaJIT.
@@
4.2.1. Just-in-Time (JIT) Warmup
JIT program execution is typically split into a startup phase and a delayed steady state phase of peak performance. As a JIT program executes, it determines which code paths are frequently executed, and compiles them into machine code. As the program continues to revisit these hotspots, performance of further iterations improves. This system allows for a good startup time to peak performance tradeoff. However, existing research has already shown that a steady state of peak performance is not always reached (Barrett et al., 2017; Traini et al., 2023). In fact, one of these studies (Barrett et al., 2017) uses the same CLBG benchmark suite and demonstrates that for some benchmarks, a steady state is never reached.
@@ 4.2.1. Just-in-Time (JIT) Warmup
Figure 8. Measuring just the first iteration is not indicative of Java’s (OpenJDK) performance for long-lived applications (§4.2.1). We gradually increase the number of iterations, observing decreased per-iteration execution time on several benchmarks, including (a) mandelbrot and (b) fannkuch-redux shown above. This effect could be due to startup and cleanup time and/or a JIT warmup effect. The curve flattens as we average over more iterations.
[…]
Two scatter plots, with x-axis “Number of iterations” and y-axis “Time per iteration [ms]”. Time per iteration decreases as the number of iterations increases, plateauing after roughly 10 iterations on each graph.
@@
Factors other than JIT warmup may also be affecting the very first iteration of those short-lived benchmarks, such as bytecode interpretation or memory and garbage collection initialization. To test this, we wrap each Java benchmark in a loop and gradually increase the number of in-process iterations. Figure 8 shows that for the mandelbrot and fannkuch-redux benchmarks, the first iteration’s execution time is 2.5×2.5$\times$2.5 × and 1.5×1.5$\times$1.5 × slower respectively than when averaging over 15 iterations. Generally, we see that for these and other benchmarks, averaging over ten or more iterations amortizes the first iteration’s impact. Applying this method over every benchmark, we observe 80% improved energy efficiency and execution time versus using only the first iteration.
4.2.2. Garbage Collection
@@
An important property of higher-level languages and their implementations is garbage collection, which imposes significant performance overhead (Hertz and Berger, 2005). Out of the 13 language implementations discussed in this paper, 10 use a garbage collector. Garbage collectors can typically be configured and tuned for each application to offer maximal performance, which is standard practice in real-world applications to optimize for application throughput and/or latency. For example, Go’s garbage collector can be disabled entirely with the GOGC environment variable. When doing so, we observe a 2.8×2.8$\times$2.8 × speedup on the allocation-intensive binary-trees benchmark.
4.3. Decomposing Energy into Power and Time
@@
Figure 9. Energy consumption is the product of power and time (§4.3). Splitting into these two components allows separate consideration of each factor. As Figure 10 shows, there is no significant difference in power draw between implementations when we keep external factors constant, hence the impact of programming language implementation on energy consumption beyond execution time is negligible.
[…]
Our previous working causal diagram, adding “Power” and “Time”.
Figure 10. Average power draw (package (PKG) + DRAM) does not differ significantly across programming language implementations (§4.3). Each point represents a single benchmark run. Points exhibit extremely low variance, and so are overlapped in the graph. After controlling for external factors such as varying number of active cores and execution time, average power draw is constant across benchmark and programming languages implementations. The red line indicates the mean power draw of $189.8\pm 0.5$ W.
[…]
A scatter plot, with x-axis “Programming Language Implementation” and y-axis “Average power draw (PKG + DRAM) [W]”. All points are essentially on the same y-value, the red line representing the mean goes through all of them.
@@
Fundamentally, energy consumption is the product of power and time. Figure 9 decomposes these two factors. To test if the choice of programming language implementation has an effect on power draw beyond execution time, we fix other factors that may affect power draw. We normalize by execution time, and exclusively for this analysis, limit the program to a single core via Docker’s --cpuset-cpus argument. To ensure frequency scaling and throttling do not increase power draw variance in our measurements, we pin CPUs to their minimum frequency with turbo mode disabled. Figure 10 shows that there is no significant difference in power draw between programming language implementations once these outside factors are controlled for.
@@
In theory, any language implementation with a runtime system running in parallel may affect energy consumption beyond variance in execution time, since increased core usage will increase power draw, as Section 4.4.1 details below. This hidden increased parallelism could occur with a parallel garbage collector or just-in-time compilation threads executing alongside the main program. In practice, compilation threads are primarily active only during startup, and both are rarely significant compared to the main program’s execution and other general overhead costs. For this reason, we draw this edge in gray in our model.
4.4. Contributors to Power Draw
@@
Figure 11. Number of active cores is the primary factor in increased power draw (§4.4). Memory activity also increases power draw, but to a much lesser extent. Figure 12 quantifies this.
[…]
Our previous working causal diagram, adding “Number of Active Cores” and “Memory Activity”.
@@
(a) Average package (PKG) power draw by average number of active cores (§4.4.1). Log regression: $y=31\log_{2}x+246\left(R^{2}=0.72\right)$.
[…]
A scatter plot with a log regression line, with x-axis “Average number of active cores” and y-axis “Average power draw (PKG) [W]”.
@@
(b) Average DRAM power draw by memory activity (§4.4.2), approximated using LLC misses. Linear regression: $y=1.68\cdot 10^{-8}x+12\left(R^{2}=0.93\right)$.
[…]
A scatter plot with a linear regression line, with x-axis “LLC misses per second [1/s]” and y-axis “Average power draw (DRAM) [W]”.
@@
Figure 12. Power draw is much more significantly affected by the number of active cores than memory activity (§4.4). Each point on those graphs represents a single benchmark run.
@@
Previous research has demonstrated that processor and memory energy usage are the main variable contributors to power draw (Vogelsang, 2010). We add these two factors to our model in Figure 11. For the studied programs on our experimental platform, the ratio of memory energy consumption to CPU energy consumption remains between 2 and 8%. This percentage is in line with the reported numbers in Pereira et al.
4.4.1. Number of Active Cores

### v3

@@ III-B2 Performance Counters
Section IV-D2 highlights the impact of memory activity on energy consumption: a majority of memory energy consumption is dependent on the rate of read and write operations. We monitor the last level cache (LLC) using performance counters to estimate memory activity. Each LLC hit means that the cache already had a copy of the data, and therefore no further memory activity occurs. On the other hand, each LLC miss means that the system must fetch data from memory. Therefore, the number of LLC misses is a good proxy for memory activity.
@@
III-C Benchmarks
This paper primarily focuses its evaluation on the benchmarks used in the original Pereira et al. study, namely the Computer Language Benchmark Game. In some analyses, we extend that evaluation with two additional benchmark suites: SPEC CPU 2017 [30] and DaCapo Chopin [31]. The following subsections provide a brief overview of each of these benchmark suites.
III-C1 Computer Language Benchmark Game
TABLE III: Descriptions of benchmarks from the Computer Language Benchmark Game (§III-C1), along with the range of lines of code (LoC) across languages for each benchmark as reported by cloc.
@@ III-C1 Computer Language Benchmark Game
binary-trees 36–98 Memory allocator stresstest
fannkuch-redux 40–158 Permutation algorithm
fasta 84–287 Random DNA generation
k-nucleotide 57–202 Frequency counting
mandelbrot 32–140 Fractal rendering
n-body 78–157 Physics simulation
pidigits 41–149 Arbitrary precision
regex-redux 22–111 String processing
reverse-complement 34–257 String transformation
spectral-norm 33–156 Linear algebra
The Computer Language Benchmark Game [20] (CLBG) is a collection of small programs implemented in many different programming languages (22 to 287 lines of code). We use the same benchmark implementations as Pereira et al. [27], which is in effect a snapshot of the fastest versions of the benchmarks as available on the CLBG repository at the time of the original work. Table III provides a brief description of each benchmark. We limit our analysis to eleven languages across thirteen implementations: C, C#, C++, Go, Java, JavaScript, Lua (Lua and LuaJIT), PHP, Python (CPython and PyPy), Rust, and TypeScript. These languages comprise 11 of the top 12 most popular general programming languages in the 2024 Stack Overflow survey [32] (excluding Kotlin).
@@
We stress that while the CLBG benchmarks are not necessarily representative of real-world applications, the causal analysis this paper develops is largely independent of the details of the benchmark implementations. It instead highlights the impact of high-level properties of the benchmark implementations, such as their degree of parallelism and cache activity.
III-C2 SPEC CPU 2017
TABLE IV: SPEC CPU 2017 benchmarks (§III-C2).
@@ III-C2 SPEC CPU 2017
perlbench C Perl interpreter
gcc C GNU C compiler
mcf C Route planning
omnetpp C++ Discrete-event simulation
xalancbmk C++ XML to HTML conversion
x264 C Video compression
deepsjeng C++ Alpha-beta tree search
leela C++ Monte Carlo tree search
exchange2 Fortran Recursive solution generator
xz C General data compression
bwaves Fortran Explosion modeling
cactuBSSN C++, C, Fortran Physics: relativity
lbm C Fluid dynamics
wrf Fortran, C Weather forecasting
cam4 Fortran, C Atmosphere modeling
pop2 Fortran, C Wide-scale ocean modeling
imagick C Image manipulation
nab C Molecular dynamics
fotonik3d Fortran Computational electromagnetics
roms Fortran Regional ocean modeling
TABLE V: DaCapo benchmarks (§III-C3).
@@
SPEC CPU 2017 [30] is a widely used benchmark suite for evaluating systems on C, C++ and Fortran code. The suite consists of large real-world programs and workloads, ranging between 1k and 1.3M lines of code. We use the latest available version (1.1.9) and run all benchmarks in the “speed” category using the reference input set. We omit a single benchmark, cam4, as it did not successfully run on our platform. Table IV briefly describes of each benchmark.
III-C3 DaCapo
DaCapo [31] is a widely used benchmark suite for Java applications. The suite consists of large real-world programs and workloads, ranging between 25k and 6M lines of code. We use the latest available version (Chopin) and run all benchmarks using the default input set. Table V provides a brief description of each benchmark.
IV Causal Model for Energy Consumption in Programming Languages
This section builds piece by piece the causal model presented as a diagram in its final form in Figure 1. Causal diagrams illustrate how different factors influence each other, with arrows indicating cause-to-effect relationships [19]. We additionally use gray arrows to represent comparatively weaker relationships, as detailed in the following sections. Each subsection below gradually adds nodes and edges to the diagram in order to obtain a rich causal model that captures the essential factors in the relationship between the choice of programming language and energy consumption. Figures 3, 4, 7, and 8 build upon each other, with new elements in orange. They culminate in the final model represented in Figure 1.
IV-A Starting Point
ProgrammingLanguageEnergy
Figure 3: The simple model used as a starting point in this paper (§IV-A). This simple model captures the relationship implied in Pereira et al., namely that the choice of programming language has a direct impact on total energy consumption.
@@ IV-A Starting Point
IV-B Programming Languages and Implementations
ProgrammingLanguageProgrammingLanguageImplementationEnergy
Figure 4: Programming languages may have multiple implementations (§IV-B), and implementation decisions such as garbage collection or using just-in-time compilation have a more direct impact on performance. We are in fact comparing implementations of programming languages, not the languages themselves.
Figure 4 introduces the important distinction between the programming languages themselves and their implementations. Some languages have a single well-known implementation, or a blurry line between language and implementation, such as Rust or Go. Others have multiple implementations, such as C/C++ with GCC, Clang, or MSVC, Java with various JVMs, Python with CPython and PyPy, or Lua with LuaJIT. Finally, there can be significant differences between specific versions of the same implementation. Our model treats those different versions as different implementations altogether. For example, recent CPython releases each successively delivered significant performance improvements, notably with the 3.11 release achieving a 25% runtime speedup over 3.10 [33].
@@ IV-B Programming Languages and Implementations
Figure 5: Programming Languages can exhibit different performance characteristics depending on the implementation used (§IV-B). Here, we compare interpreters (CPython, Lua) to their JIT equivalent (PyPy, LuaJIT) on the same CLBG benchmark sources. We normalize to the interpreter results. We find around 5×5$\times$5 × and 1.25×1.25$\times$1.25 × decreased execution time and energy consumption on average for LuaJIT over Lua and PyPy over CPython, respectively. These results demonstrate the impact of choice of programming language implementation versus the language itself.
@@
IV-B1 Just-in-Time (JIT) Warmup
JIT program execution is typically split into a startup phase and a delayed steady state phase of peak performance. As a JIT program executes, it determines which code paths are frequently executed, and compiles them into machine code. As the program continues to revisit these hotspots, performance of further iterations improves. This system allows for a good startup time to peak performance tradeoff. However, existing research has already shown that a steady state of peak performance is not always reached [24, 23]. In fact, one of these studies [24] uses the same CLBG benchmark suite and demonstrates that for some benchmarks, a steady state is never reached.
@@ IV-B1 Just-in-Time (JIT) Warmup
Figure 6: Measuring just the first iteration is not indicative of Java’s (OpenJDK) performance for long-lived applications (§IV-B1). We gradually increase the number of iterations, observing decreased per-iteration execution time on several benchmarks, including (a) mandelbrot and (b) fannkuch-redux shown above. This effect could be due to startup and cleanup time and/or a JIT warmup effect. The curve flattens as we average over more iterations.
@@
Factors other than JIT warmup may also affect the first iteration of short-lived benchmarks, such as bytecode interpretation or memory and garbage collection initialization. To verify this, we wrap Java benchmarks in a loop and gradually increase the number of in-process iterations. Figure 6 shows that for the mandelbrot and fannkuch-redux benchmarks, the first iteration’s execution time is respectively 2.7×2.7$\times$2.7 × and 1.4×1.4$\times$1.4 × slower than when averaging over 15 iterations. Generally, we see that for these and other benchmarks, averaging over ten or more iterations amortizes the first iteration’s impact: across the CLBG benchmarks, we observe a mean of 80% improved energy efficiency and execution time.
IV-B2 Garbage Collection
@@
An important property of higher-level languages and their implementations is garbage collection, which imposes significant performance overhead [25]. Out of the 13 language implementations discussed in this paper, 10 use a garbage collector. Garbage collectors can typically be configured and tuned for each application to offer maximal performance, which is standard practice in real-world applications to optimize for application throughput and/or latency. For example, Go’s garbage collector can be disabled entirely with the GOGC environment variable. When doing so, we observe a 2.8×2.8$\times$2.8 × speedup on the allocation-intensive binary-trees benchmark, and a 1.23×1.23$\times$1.23 × speedup on mandelbrot.
IV-C Decomposing Energy into Power and Time
ProgrammingLanguageProgrammingLanguageImplementationTimePowerEnergy
@@ IV-C Decomposing Energy into Power and Time
Figure 7: Energy consumption is the product of power and time (§IV-C). Splitting into these two components allows separate consideration of each factor.
@@
Fundamentally, energy consumption is the product of power and time. Figure 7 decomposes these two factors. To test if the choice of programming language implementation has an effect on power draw beyond execution time, we fix other factors that may affect power draw: we limit the program to a single core via Docker’s --cpuset-cpus argument, and pin CPUs to their minimum frequency with turbo mode disabled to ensure frequency scaling and throttling do not increase power draw variance in our measurements. Once these factors are controlled for, we measure runtime and energy consumption of each benchmark across all language implementations. These measurements yield equal power draw: $189.8\pm 0.5$ W. Hence, the impact of choice of programming language implementation (or programming language more generally) on energy consumption beyond execution time is negligible.
@@
In theory, any language implementation with a runtime system running in parallel may affect energy consumption beyond variance in execution time, since increased core usage will increase power draw, as Section IV-D1 details below. This hidden increased parallelism could occur with a parallel garbage collector or just-in-time compilation threads executing alongside the main program. In practice, compilation threads are primarily active only during startup, and both are rarely significant compared to the main program’s execution and other general overhead costs. For this reason, we draw this edge in gray in our model.
IV-D Contributors to Power Draw
ProgrammingLanguageProgrammingLanguageImplementationTimeNumber ofActive CoresMemoryActivityPowerEnergy
@@ IV-D Contributors to Power Draw
Figure 8: Number of active cores is the primary factor in increased power draw (§IV-D). Memory activity also increases power draw, but to a much lesser extent. Figure 9 quantifies this.
@@
(a) Average package (PKG) power draw by average number of active cores (§IV-D1). Log regression: $y=31\log_{2}x+246\left(R^{2}=0.72\right)$.
@@
(b) Average DRAM power draw by memory activity (§IV-D2), approximated using LLC misses. Linear regression: $y=1.68\cdot 10^{-8}x+12\left(R^{2}=0.93\right)$.
@@
Figure 9: Power draw is much more significantly affected by the number of active cores than memory activity (§IV-D). Each point on those graphs represents a single benchmark run.
@@
Previous research has demonstrated that processor and memory energy usage are the main variable contributors to power draw [34]. We add these two factors to our model in Figure 8. For the studied programs on our experimental platform, the ratio of memory energy consumption to CPU energy consumption remains between 2 and 8%. This percentage is in line with the reported numbers in Pereira et al.
IV-D1 Number of Active Cores